### v1

@@ 1 Introduction
In recent years, optimal transport-based techniques like the Wasserstein distance and variations of this have become popular tools for image comparison. At their core, the Wasserstein distances define metrics between probability measures on a common Polish space and find an optimal coupling [1, 2, 3, 4]. Although being based on a linear program, the numerical calculation is challenging, initiating the development of entropic optimal transport [5], linear optimal transport [6, 7, 8], and the so-called sliced Wasserstein distance [9, 10, 11, 12] to lower the computational burden. The definition of the Wasserstein distance can be extended to ensure invariance under orthogonal transformations of the input measures, leading to the so-called Procrustes–Wasserstein distance [13, 14].
@@
  • a rigours discussion about the linear separability properties, which provably carry over to all new extended NR-CDT variants,
@@ 2.1 Radon Transform of Functions
For $f\in L^{1}(\mathbb{R}^{2})$, we define its Radon transform $\mathcal{R}[f]\colon\mathbb{R}\times\mathbb{S}_{1}\to\mathbb{R}$ as the line integral
@@
where $\mathop{}\!\mathrm{d}s$ denotes the arc length element of $\ell_{t,{\bm{\theta}}}$. This defines the Radon operator $\mathcal{R}\colon L^{1}(\mathbb{R}^{2})\to L^{1}(\mathbb{R}\times\mathbb{S}_{1})$. For fixed ${\bm{\theta}}\in\mathbb{S}_{1}$, we set $\mathcal{R}_{\bm{\theta}}\coloneqq\mathcal{R}(\cdot,{\bm{\theta}})$, which is referred to as the restricted Radon operator $\mathcal{R}_{\bm{\theta}}\colon L^{1}(\mathbb{R}^{2})\to L^{1}(\mathbb{R})$. The action of the Radon operator is illustrated in Figure 1.
@@
The Radon transform is also well-defined for all $f\in L^{p}(\mathbb{R}^{2})$ with $p\geq 1$ and $\operatorname{supp}(f)\subseteq\mathbb{B}_{2}\coloneqq\{\mathbf{x}\in\mathbb{R}^{2}\mid\lVert\mathbf{x}\rVert\leq 1\}$, in which case $\mathcal{R}[f]\in L^{p}(\mathbb{R}\times\mathbb{S}_{1})$ with $\operatorname{supp}(\mathcal{R}[f])\subseteq\mathbb{I}\times\mathbb{S}_{1}$, where $\mathbb{I}\coloneqq[-1,1]$. According to [25], the adjoint operator $\mathcal{R}^{*}\colon L^{\infty}(\mathbb{R}\times\mathbb{S}_{1})\to L^{\infty}(\mathbb{R}^{2})$ of the Radon transform $\mathcal{R}\colon L^{1}(\mathbb{R}^{2})\to L^{1}(\mathbb{R}\times\mathbb{S}_{1})$ is given by the back projection
@@ 2.2 Radon Transform of Measures
One can find the measure-valued Radon transform $\mathcal{R}\colon\mathcal{M}(\mathbb{R}^{2})\to\mathcal{M}(\mathbb{R}\times\mathbb{S}_{1})$ as the adjoint of the function-valued adjoint $\mathcal{R}^{*}\colon L^{\infty}(\mathbb{R}\times\mathbb{S}_{1})\to L^{\infty}(\mathbb{R}^{2})$, similar to the case of distributions with compact support, cf. [24].
@@
$\langle\mathcal{R}[\mu],g\rangle=\langle\mu,\mathcal{R}^{*}[g]\rangle\quad\forall\,g\in L^{\infty}(\mathbb{R}\times\mathbb{S}_{1}).$
@@
For all $\mu\in\mathcal{M}(\mathbb{R}^{2})$ and $g\in L^{\infty}(\mathbb{R}\times\mathbb{S}_{1})$, applying Fubini’s theorem gives
@@
Note that, for $f\in L^{1}(\mathbb{R}^{2})$ and the Lebesgue measure $\lambda_{\mathbb{R}^{2}}$ on $\mathbb{R}^{2}$, the Radon transform satisfies
@@ 3.1 R-CDT for Measures
$\widehat{\mu}\coloneqq F_{\mu}^{[-1]}\circ F_{\rho}.$
@@
$\widehat{\mu}=\operatorname*{arg\,min}_{T_{\#}\rho=\mu}\int_{\mathbb{R}}c(s-T(s))\mathop{}\!\mathrm{d}\rho(s),$
where the minimum is taken over all measurable functions $T\colon\mathbb{R}\to\mathbb{R}$. In other words, $\widehat{\mu}\colon\mathbb{R}\to\mathbb{R}$ is an optimal Monge map transporting $\rho$ to $\mu$ while minimizing the cost. If $\mu\in\mathcal{P}_{2}(\mathbb{R})$, i.e., $\mu$ has finite 2nd moment, then $\widehat{\mu}$ is square integrable with respect to $\rho$, i.e., $\widehat{\mu}\in L^{2}_{\rho}(\mathbb{R})$. Moreover, for $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R})$, the norm distance
@@
If $\mu\in\mathcal{P}_{2}(\mathbb{R}^{2})$, then the Radon projection $\mathcal{R}_{\bm{\theta}}[\mu]\in\mathcal{P}_{2}(\mathbb{R})$ has finite 2nd moment as well. Consequently, $\widehat{\mathcal{R}}[\mu]\in L_{\rho\times u_{\mathbb{S}_{1}}}^{2}(\mathbb{R}\times\mathbb{S}_{1})$. For $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{2})$, the norm distance
@@ 3.2 Normalized R-CDT
Here, $\subset\joinrel\subset$ denotes a compact subset, and $\dim$ the dimension of the affine hull. For these, the standard deviation of the restricted Radon transform is bounded away from zero and cannot vanish.
@@ 3.3 Max-Normalized R-CDT
We show that $\mathcal{N}_{\mathrm{m}}$ maps a given measure to a bounded function so that the $\mathrm{m}$NR-CDT space $\mathcal{N}_{\mathrm{m}}[\mathcal{P}_{\mathrm{c}}^{*}(\mathbb{R}^{2})]$ is contained in $L_{\rho}^{\infty}(\mathbb{R})$ for the underlying reference measure $\rho\in\mathcal{P}(\mathbb{R})$.
@@
Let $\mu\in\mathcal{P}_{\mathrm{c}}^{*}(\mathbb{R}^{2})$. Then, $\mathcal{N}_{\mathrm{m}}[\mu]\in L^{\infty}_{\rho}(\mathbb{R})$.
@@
Due to the affine construction of $\mathbb{F}$ and $\mathbb{G}$, Proposition 7 yields $\mathcal{N}_{\mathrm{m}}[\mathbb{F}]=\bigl\{\mathcal{N}_{\mathrm{m}}[\mu_{0}]\bigr\}$ and $\mathcal{N}_{\mathrm{m}}[\mathbb{G}]=\bigl\{\mathcal{N}_{\mathrm{m}}[\nu_{0}]\bigr\}$. Hence, the assumption $\mathcal{N}_{\mathrm{m}}[\mu_{0}]\neq\mathcal{N}_{\mathrm{m}}[\nu_{0}]$ implies the linear separability of $\mathcal{N}_{\mathrm{m}}[\mathbb{F}]$ and $\mathcal{N}_{\mathrm{m}}[\mathbb{G}]$ in $L^{\infty}_{\rho}(\mathbb{R})$. ∎
The proof of Theorem 1 reveals that the classes $\mathbb{F}$ and $\mathbb{G}$ collapse to single points in $\mathrm{m}$NR-CDT space, directly allowing for linear separability. This is in contrast to the R-CDT results in [23], which can only tackle a subset of affine transforms, and where the considered classes form convex sets in R-CDT space.
@@ 3.4 $h$-Normalized R-CDT
We now describe a more general final normalization step than proposed in § 3.3 for treating the resorting of $\{\mathcal{N}_{\bm{\theta}}[\mu]\mid{\bm{\theta}}\in\mathbb{S}_{1}\}$. To this end, let $h\colon L^{\infty}(\mathbb{S}_{1})\to\mathbb{R}$ be boundedness-preserving, i.e., bounded sets are mapped to bounded sets, such that $h(g\circ\phi_{\mathbf{A}})=h(g)$ for all $g\in\mathrm{L}^{\infty}(\mathbb{S}_{1})$ and $\phi_{\mathbf{A}}({\bm{\theta}})=(\mathbf{A}^{\top}{\bm{\theta}})/\lVert\mathbf{A}^{\top}{\bm{\theta}}\rVert$ with $\mathbf{A}\in\mathrm{GL}(2)$. With this, for $\mu\in\mathcal{P}_{\mathrm{c}}^{*}(\mathbb{R}^{2})$, we define its $h$-normalized R-CDT ($h_{\mathrm{}}$NR-CDT) $\mathcal{N}_{h_{\mathrm{}}}[\mu]\colon\mathbb{R}\to\mathbb{R}$ via
@@
We show that $\mathcal{N}_{h_{\mathrm{}}}$ maps a given measure to a bounded function so that the $h_{\mathrm{}}$NR-CDT space $\mathcal{N}_{h_{\mathrm{}}}[\mathcal{P}_{\mathrm{c}}^{*}(\mathbb{R}^{2})]$ is contained in $L_{\rho}^{\infty}(\mathbb{R})$ for the underlying reference measure $\rho\in\mathcal{P}(\mathbb{R})$.
@@
$\mathcal{N}_{h_{\mathrm{}}}[\mu]\in L^{\infty}_{\rho}(\mathbb{R})$ for all $\mu\in\mathcal{P}_{\mathrm{c}}^{*}(\mathbb{R}^{2})$.
@@
and $c\coloneqq\min_{{\bm{\theta}}\in\mathbb{S}_{1}}\operatorname{std}(\widehat{\mathcal{R}}_{\bm{\theta}}[\mu])>0$. Thus, the NR-CDT is uniformly bounded by $\lvert\mathcal{N}[\mu](t,{\bm{\theta}})\rvert\leq\operatorname{diam}(\mu)/c$ for all $t\in\mathbb{R}$ and ${\bm{\theta}}\in\mathbb{S}_{1}$ so that $\mathcal{N}[\mu](t,\cdot)\in L^{\infty}(\mathbb{S}_{1})$. Since $h\colon L^{\infty}(\mathbb{S}_{1})\to\mathbb{R}$ is bounded, there exists a constant $C>0$ such that $\lvert\mathcal{N}_{h_{\mathrm{}}}[\mu](t)\rvert\leq C$ for all $t\in\mathbb{R}$. ∎
@@
By construction of $\mathbb{F}$ and $\mathbb{G}$, Proposition 9 yields $\mathcal{N}_{h_{\mathrm{}}}[\mathbb{F}]=\bigl\{\mathcal{N}_{h_{\mathrm{}}}[\mu_{0}]\bigr\}$ and $\mathcal{N}_{h_{\mathrm{}}}[\mathbb{G}]=\bigl\{\mathcal{N}_{h_{\mathrm{}}}[\nu_{0}]\bigr\}$. Hence, the assumption $\mathcal{N}_{h_{\mathrm{}}}[\mu_{0}]\neq\mathcal{N}_{h_{\mathrm{}}}[\nu_{0}]$ implies the linear separability of $\mathcal{N}_{h_{\mathrm{}}}[\mathbb{F}]$ and $\mathcal{N}_{h_{\mathrm{}}}[\mathbb{G}]$ in $L^{\infty}_{\rho}(\mathbb{R})$. ∎
@@
$h(g)=H(\{\eta(g({\bm{\theta}}))\mid{\bm{\theta}}\in\mathbb{S}_{1}\}),\quad g\in L^{\infty}(\mathbb{S}_{1}),$

### v2

@@ 1 Introduction
In recent years, optimal transport-based techniques like the Wasserstein distance and variations of this have become popular tools for image comparison. At their core, the Wasserstein distances define metrics between probability measures on a common Polish space and find an optimal coupling [1, 2, 3, 4]. Although being based on a linear program, the numerical calculation is challenging, initiating the development of entropic optimal transport [5, 6], linear optimal transport [7, 8, 9], so-called sliced Wasserstein distance [10, 11, 12, 13, 14], and the subspace robust Wasserstein distance [15] to lower the computational burden and to increase stability. The definition of the Wasserstein distance can be extended to ensure invariance under orthogonal transformations of the input measures, leading to the so-called Procrustes–Wasserstein distance [16, 17] and the rotation-invariant sliced Wasserstein distance [18, 14].
@@
  • a rigorous discussion about the linear separability properties, which provably carry over to all new extended NR-CDT variants,
@@ 2.1 Radon Transform of Functions
For $f\in\mathrm{L}^{1}(\mathbb{R}^{2})$, we define its Radon transform $\mathcal{R}[f]\colon\mathbb{R}\times\mathbb{S}_{1}\to\mathbb{R}$ as the line integral
@@
where $\mathop{}\!\mathrm{d}s$ denotes the arc length element of $\ell_{t,{\bm{\theta}}}$. This defines the Radon operator $\mathcal{R}\colon\mathrm{L}^{1}(\mathbb{R}^{2})\to\mathrm{L}^{1}(\mathbb{R}\times\mathbb{S}_{1})$. For fixed ${\bm{\theta}}\in\mathbb{S}_{1}$, we set $\mathcal{R}_{\bm{\theta}}\coloneqq\mathcal{R}(\cdot,{\bm{\theta}})$, which is referred to as the restricted Radon operator $\mathcal{R}_{\bm{\theta}}\colon\mathrm{L}^{1}(\mathbb{R}^{2})\to\mathrm{L}^{1}(\mathbb{R})$. The action of the Radon operator is illustrated in Figure 1.
@@
The Radon transform is also well-defined for all $f\in\mathrm{L}^{p}(\mathbb{R}^{2})$ with $p\geq 1$ and $\operatorname{supp}(f)\subseteq\mathbb{B}_{2}\coloneqq\{\mathbf{x}\in\mathbb{R}^{2}\mid\lVert\mathbf{x}\rVert\leq 1\}$, in which case $\mathcal{R}[f]\in\mathrm{L}^{p}(\mathbb{R}\times\mathbb{S}_{1})$ with $\operatorname{supp}(\mathcal{R}[f])\subseteq\mathrm{i}\times\mathbb{S}_{1}$, where $\mathrm{i}\coloneqq[-1,1]$. According to [29], the adjoint operator $\mathcal{R}^{*}\colon\mathrm{L}^{\infty}(\mathbb{R}\times\mathbb{S}_{1})\to\mathrm{L}^{\infty}(\mathbb{R}^{2})$ of the Radon transform $\mathcal{R}\colon\mathrm{L}^{1}(\mathbb{R}^{2})\to\mathrm{L}^{1}(\mathbb{R}\times\mathbb{S}_{1})$ is given by the back projection
@@ 2.2 Radon Transform of Measures
One can find the measure-valued Radon transform $\mathcal{R}\colon\mathcal{M}(\mathbb{R}^{2})\to\mathcal{M}(\mathbb{R}\times\mathbb{S}_{1})$ as the adjoint of the function-valued adjoint $\mathcal{R}^{*}\colon\mathrm{L}^{\infty}(\mathbb{R}\times\mathbb{S}_{1})\to\mathrm{L}^{\infty}(\mathbb{R}^{2})$, similar to the case of distributions with compact support, cf. [28].
@@
$\langle\mathcal{R}[\mu],g\rangle=\langle\mu,\mathcal{R}^{*}[g]\rangle\quad\forall\,g\in\mathrm{L}^{\infty}(\mathbb{R}\times\mathbb{S}_{1}).$
@@
For all $\mu\in\mathcal{M}(\mathbb{R}^{2})$ and $g\in\mathrm{L}^{\infty}(\mathbb{R}\times\mathbb{S}_{1})$, applying Fubini’s theorem gives
@@
Note that, for $f\in\mathrm{L}^{1}(\mathbb{R}^{2})$ and the Lebesgue measure $\lambda_{\mathbb{R}^{2}}$ on $\mathbb{R}^{2}$, the Radon transform satisfies
@@ 3.1 R-CDT for Measures
$\widehat{\mu}\coloneqq F_{\mu}^{[-1]}\circ F_{\rho}.$ (3)
Throughout the paper we use a fixed $\rho$, but all our results are independent from it’s choice. For $\rho=u_{[0,1]}$, the CDT reduces to $\widehat{\mu}=F_{\mu}^{[-1]}$. In applications, however, a different choice of $\rho$ might be advantageous.
@@
$\widehat{\mu}=\operatorname*{arg\,min}_{T_{\#}\rho=\mu}\int_{\mathbb{R}}c(s-T(s))\mathop{}\!\mathrm{d}\rho(s),$ (4)
where the minimum is taken over all measurable functions $T\colon\mathbb{R}\to\mathbb{R}$. In other words, $\widehat{\mu}\colon\mathbb{R}\to\mathbb{R}$ is an optimal Monge map transporting $\rho$ to $\mu$ while minimizing the cost. If $\mu\in\mathcal{P}_{2}(\mathbb{R})$, i.e., $\mu$ has finite 2nd moment, then $\widehat{\mu}$ is square integrable with respect to $\rho$, i.e., $\widehat{\mu}\in\mathrm{L}^{2}_{\rho}(\mathbb{R})$. Moreover, for $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R})$, the norm distance
@@
If $\mu\in\mathcal{P}_{2}(\mathbb{R}^{2})$, then the Radon projection $\mathcal{R}_{\bm{\theta}}[\mu]\in\mathcal{P}_{2}(\mathbb{R})$ has finite 2nd moment as well. Consequently, $\widehat{\mathcal{R}}[\mu]\in\mathrm{L}_{\rho\times u_{\mathbb{S}_{1}}}^{2}(\mathbb{R}\times\mathbb{S}_{1})$. For $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{2})$, the norm distance
@@
Another generalization of the CDT to multi-dimensional domains is the so-called linear optimal transport (LOT) [9]. This approach is based on the fact that the CDT computes an optimal Monge map, i.e., a solution to (4). Similarly, LOT maps a given measure to an optimal multi-dimensional Monge map $T\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$. The key feature of both transforms—R-CDT and LOT—is that they allow for linear separation of certain measure classes. The theoretical advantage of LOT is that this mapping can be interpreted as lifting to the Wasserstein tangent space, which is not the case for R-CDT. On the downside, the computation of LOT is numerical more challenging and requires approximations. In contrast to this, the R-CDT can be computed analytically.
@@ 3.2 Normalized R-CDT
Here, $\subset\joinrel\subset$ denotes a compact subset, and $\dim$ the dimension of the affine hull. For these, the standard deviation of the restricted Radon transform is bounded away from zero and cannot vanish. If $\dim(\operatorname{supp}(\mu))=1$, $\mu$ is supported on a line and $\mathcal{N}_{\bm{\theta}}[\mu]$ is well-defined for all ${\bm{\theta}}\in\mathbb{S}^{1}$ except the two being orthogonal to the line. If $\mu$ is a single point measure, $\mathcal{N}_{\bm{\theta}}[\mu]$ is not well-defined for any ${\bm{\theta}}\in\mathbb{S}^{1}$.
@@ 3.3 Max-Normalized R-CDT
We show that $\mathcal{N}_{\mathrm{m}}$ maps a given measure to a bounded function so that the $\mathrm{m}$NR-CDT space $\mathcal{N}_{\mathrm{m}}[\mathcal{P}_{\mathrm{c}}^{*}(\mathbb{R}^{2})]$ is contained in $\mathrm{L}_{\rho}^{\infty}(\mathbb{R})$ for the underlying reference measure $\rho\in\mathcal{P}(\mathbb{R})$.
@@
Let $\mu\in\mathcal{P}_{\mathrm{c}}^{*}(\mathbb{R}^{2})$. Then, $\mathcal{N}_{\mathrm{m}}[\mu]\in\mathrm{L}^{\infty}_{\rho}(\mathbb{R})$.
@@
Due to the affine construction of $\mathbb{F}$ and $\mathbb{G}$, Proposition 7 yields $\mathcal{N}_{\mathrm{m}}[\mathbb{F}]=\bigl\{\mathcal{N}_{\mathrm{m}}[\mu_{0}]\bigr\}$ and $\mathcal{N}_{\mathrm{m}}[\mathbb{G}]=\bigl\{\mathcal{N}_{\mathrm{m}}[\nu_{0}]\bigr\}$. Hence, the assumption $\mathcal{N}_{\mathrm{m}}[\mu_{0}]\neq\mathcal{N}_{\mathrm{m}}[\nu_{0}]$ implies the linear separability of $\mathcal{N}_{\mathrm{m}}[\mathbb{F}]$ and $\mathcal{N}_{\mathrm{m}}[\mathbb{G}]$ in $\mathrm{L}^{\infty}_{\rho}(\mathbb{R})$. ∎
The proof of Theorem 1 reveals that the classes $\mathbb{F}$ and $\mathbb{G}$ collapse to single points in $\mathrm{m}$NR-CDT space, directly allowing for linear separability. This theorem and the following generalizations are stated in this form to highlight the analogy to R-CDT [27] and LOT [9, 8], where linear separability is observed as well as one of the key achievements. The transformation of affine classes to singletons, however, is in contrast to the results in the literature, which can only tackle a subset of affine transforms, and where the considered classes are mapped to convex sets.
@@ 3.4 $h$-Normalized R-CDT
We now describe a more general final normalization step than proposed in § 3.3 for treating the resorting of $\{\mathcal{N}_{\bm{\theta}}[\mu]\mid{\bm{\theta}}\in\mathbb{S}_{1}\}$. To this end, let the aggregation function $h\colon\mathrm{L}^{\infty}(\mathbb{S}_{1})\to\mathbb{R}$ be boundedness-preserving, i.e., bounded sets are mapped to bounded sets, such that $h(g\circ\phi_{\mathbf{A}})=h(g)$ for all $g\in\mathrm{L}^{\infty}(\mathbb{S}_{1})$ and $\phi_{\mathbf{A}}({\bm{\theta}})=(\mathbf{A}^{\top}{\bm{\theta}})/\lVert\mathbf{A}^{\top}{\bm{\theta}}\rVert$ with $\mathbf{A}\in\mathrm{GL}(2)$. With this, for $\mu\in\mathcal{P}_{\mathrm{c}}^{*}(\mathbb{R}^{2})$, we define its $h$-normalized R-CDT ($h_{\mathrm{}}$NR-CDT) $\mathcal{N}_{h_{\mathrm{}}}[\mu]\colon\mathbb{R}\to\mathbb{R}$ via
@@
We show that $\mathcal{N}_{h_{\mathrm{}}}$ maps a given measure to a bounded function so that the $h_{\mathrm{}}$NR-CDT space $\mathcal{N}_{h_{\mathrm{}}}[\mathcal{P}_{\mathrm{c}}^{*}(\mathbb{R}^{2})]$ is contained in $\mathrm{L}_{\rho}^{\infty}(\mathbb{R})$ for the underlying reference measure $\rho\in\mathcal{P}(\mathbb{R})$.
@@
$\mathcal{N}_{h_{\mathrm{}}}[\mu]\in\mathrm{L}^{\infty}_{\rho}(\mathbb{R})$ for all $\mu\in\mathcal{P}_{\mathrm{c}}^{*}(\mathbb{R}^{2})$.
@@
and $c\coloneqq\min_{{\bm{\theta}}\in\mathbb{S}_{1}}\operatorname{std}(\widehat{\mathcal{R}}_{\bm{\theta}}[\mu])>0$. Thus, the NR-CDT is uniformly bounded by $\lvert\mathcal{N}[\mu](t,{\bm{\theta}})\rvert\leq\operatorname{diam}(\mu)/c$ for all $t\in\mathbb{R}$ and ${\bm{\theta}}\in\mathbb{S}_{1}$ so that $\mathcal{N}[\mu](t,\cdot)\in\mathrm{L}^{\infty}(\mathbb{S}_{1})$. Since $h\colon\mathrm{L}^{\infty}(\mathbb{S}_{1})\to\mathbb{R}$ is bounded, there exists a constant $C>0$ such that $\lvert\mathcal{N}_{h_{\mathrm{}}}[\mu](t)\rvert\leq C$ for all $t\in\mathbb{R}$. ∎
@@
By construction of $\mathbb{F}$ and $\mathbb{G}$, Proposition 9 yields $\mathcal{N}_{h_{\mathrm{}}}[\mathbb{F}]=\bigl\{\mathcal{N}_{h_{\mathrm{}}}[\mu_{0}]\bigr\}$ and $\mathcal{N}_{h_{\mathrm{}}}[\mathbb{G}]=\bigl\{\mathcal{N}_{h_{\mathrm{}}}[\nu_{0}]\bigr\}$. Hence, the assumption $\mathcal{N}_{h_{\mathrm{}}}[\mu_{0}]\neq\mathcal{N}_{h_{\mathrm{}}}[\nu_{0}]$ implies the linear separability of $\mathcal{N}_{h_{\mathrm{}}}[\mathbb{F}]$ and $\mathcal{N}_{h_{\mathrm{}}}[\mathbb{G}]$ in $\mathrm{L}^{\infty}_{\rho}(\mathbb{R})$. ∎
@@
$h(g)=H(\{\eta(g({\bm{\theta}}))\mid{\bm{\theta}}\in\mathbb{S}_{1}\}),\quad g\in\mathrm{L}^{\infty}(\mathbb{S}_{1}),$